Ridley Inc. is a small, publicly traded construction company with 50 million shares outstanding trading at...

  1. Ridley Inc. is a small, publicly traded construction company with 50 million shares outstanding trading at $10 a share. Ridley has:
  • Unlevered beta of 1.1
  • One bank loan with interest payments of $5 million a year for 5 years and a principal payment (FV) of $250 million
  • Rating of Baa, for which the default spread is 2%
  • Marginal tax rate of 40%

The risk-free rate is 2.5%, and the equity risk premium is 5.5%. What is the current cost of capital for Ridley? (hint: find the MV of debt using the data provided … in order words find the PV of the debt)
Answer #1

Given for Ridely Inc.

number of shares outstanding = 50 million

current share price = $10

So, Market value of equity = $500 million

company has one bank loan with interest payments of $5 million a year for 5 years and a principal payment (FV) of $250 million

default spread = 2% and risk free rate rf = 2.5%

So, cost of debt Kd = Rf + default spread = 2+2.5 = 4.5%

using cost of debt as yield, current market price of debt is sum of PV of all coupon and face value discounted at Kd

So, MV of debt = 5/1.045 + 5/1.045^2 + 5/1.045^3 + 5/1.045^4 + 255/1.045^5 = $222.56 million

unlevered beta = 1.1 and tax rate T = 40%

So, levered beta = unlevered beta*(1 + (1-T)*D/E)) = 1.1*(1+0.6*222.56/500) = 1.39

equity market risk premium MRP = 5.5%

So, cost of equity using CAPM = Rf + Beta*MRP = 2.5 + 1.39*5.5 = 10.17%

current cost of capital for Ridley = We*Ke + Wd*Kd*(1-T)

We = equity/(debt+equity) = 500/(500+222.56) = 0.69

Wd = debt/(debt+equity) = 222.56/(222.56+500) = 0.31

So, WACC = 0.69*10.17 + 0.31*4.5*(1-0.4) = 7.87%

current cost of capital for Ridley = 7.87%